# Recitation over Reasoning: How Cutting-Edge Language Models Can Fail on Elementary School-Level Reasoning Problems?

### v2

@@ 4.1.2 Vision-based Problems
Evalulations. We evaluate $15$ cutting-edge VLMs, which are: GPT-4.5-Preview, OpenAI-o1-1217, GPT-4o-1120, Gemini-2.0 Pro-0205, GPT-4o-mini-0718, Gemini-2.0 Flash-0121, Qwen-2.5-VL-max, GLM-4v-Plus, Qwen-2.5-VL-72B, Claude 3.5 Sonnet, StepFun-1v-32k, Nova-Pro, Claude 3.7 Sonnet, SenseChat-Vision [59], and Qwen2.5-VL-7B. Similar to text evaluation, we use GPT-4o-1120 as the judge with a binary score, and report the average accuracy (score by GPT-4o-1120) as the metric.
@@
Table 2: Results on vision-based problems of RoR-Bench, sorted by original score accuracy. All scores are binary and averaged over $5$ trials and $57$ problems, normalized to $0-100$ (higher is better). Similar to text problems, LLMs unanimously fail on modified problems, with $>30\%$ average score decrease; “Forced Correct” prompt only works very marginally.
Results. Tab. 2 shows the result for all VLMs tested on RoR-Bench, which exhibits a $>35\%$ perfomance decrease on average from original problems to the modified problems. Interestingly, we find GPT-4o-1120, GPT-4.5-Preview and OpenAI-o1-1217 to be significantly better on original problems, but much worse on modified problems; upon checking responses, we find that the OpenAI models listed above are much more likely to summarize the origin of the images, as we collect them usually from illustrations of famous visual effects (e.g. Mach bands and checker-shadow illusions). On the contrary, models like Claude 3.5 Sonnet and Claude 3.7 Sonnet usually do not explicitly summarize such visual effects. Such result indicates that 1) OpenAI models may be overfitting to usual test cases, and more importantly, 2) explicit summarization or knowledge retrieval, which already becomes a common practice for prompt-engineering works [37, 85], is a double-edged sword; while they improve the performance on usual test cases, it may increase the risk of missing key details in the problem during summarization.
@@ 4.2 Is Few-Shot In-Context Learning the Cure?
Evaluations. We evaluate the same set of LLMs in Sec. 4.1.1  [^4] For better readability, we only show the most representative models in the main paper; see Appendix B.3 for details. For case 1 (adding original problems) mentioned above, we add a simple prompt mentioning the original problem and solution are an example (See Appendix A.3 for details). For case 2 (adding modified problems), we uniformly randomly select modified problems other than the current problem as shots; we test both 1-shot and 5-shot scenario.
Results. The results of the most represenatitive LLMs for few-shot In-Context Learning (ICL) is listed in Tab. 3; see Appendix B.3 for other LLMs. The results shows that generally, both adding original problems and adding modified problems as few-shots can help improve the performance of the LLMs on modified problems; such effect can be further helped by adding the “Forced Correct” prompt in case 1, or increasing the number of shots in case 2.
Therefore, such fixes can be seen as an initial solution; however, the performance gap between all these fixes and original problems is still very large ($>30\%$), which indicates that few-shot ICL is not the ideal panecea for LLMs to overcome the recitation issue.
@@
Table 3: The results of adding original problems as 1-shot (case 1) or adding other modified problems as few-shot (case 2) sorted by average score on modified problems in our benchmark. Claude 3.7 Sonnet and Gemini-2.0 Flash-0121 are without long CoT. Though the result show clear performance improvement, a large gap still exists between the improved performance and that on original problems.
4.3 The Mental Seal of Solvability
In Cixin Liu’s famous sci-fi novel The Dark Forest [40], the “mental seal” is a technique that injects certain statements into human brain, and can lead to firm belief of the statements even when they contradict with ground truths. As we examine the “no solution” problems in our benchmark (see Sec. 3.2 for details), we found that LLMs are particularly worse in correctly pointing out the problems with no solution, and often will make mistakes to make up a solution, as if injected by the mental seal that the problem is definitely solvable.
Evaluations. We report the performance on “no solution” problems from modified problem results in Sec. 4.1.1. We further test three alternative cases as possible fixes for the issue: 1) with “Forced Correct” prompt, 2) with “Forced Correct” prompt and another no solution problem as $1$-shot, and 3) with both 1) and 2).
@@ 4.3 The Mental Seal of Solvability
Table 4: The scores for “no solution” problems and possible fixes, sorted by average score on such of problems. Claude 3.7 Sonnet and Gemini-2.0 Flash-0121 are without long CoT. It is clearly shown that without any fixes, the average score for “no solution” problems is extremely low, showing the firm belief of LLMs that the given problem is solvable. While some LLMs, such as GPT-4.5-Preview, can be effectively corrected by adding “Forced Correct” (FC) prompts and other “no solution” problems as $1$-shot, other LLMs such as DeepSeek-R1 are still very stubborn.
Results. Tab. 4 shows the performance of the most representative LLMs on “no solution” problems as stated in Sec. 3.2 (see Appendix B.4 for other LLMs). Surprisingly, without any fixes, LLMs are unanimously stubborn on the belief that the given problem is solvable; not a single model achieves $>15\%$ score on this type of problems. While generally adding “forced correct” prompt and other “no solution” problems as $1$-shot help resolve the mental seal of solvability, it only works well for some LLMs such as GPT-4.5-Preview, and is generally still far from satisfactory for most models; for LLMs such as DeepSeek-R1 and many other weaker models, such as Qwen small models (see Tab. 10 in Appendix B.4), the issue persists.
@@ 5 Discussion and Conclusion
In this work, we propose RoR-Bench, a multimodal Chinese benchmark which clearly reveals an alarming issue in the current that current cutting-edge LLMs are unable to address even simple reasoning problems with conditions subtly shifted. Such phenomenon proved that LLMs are conducting recitation instead of reasoning when confronting seemingly classic problems. We found such issue can lead to dramatic performance loss ($>50\%$) and is unable to be addressed by simple fixes such as adding instruction prompts or few-shots, indicating that such issue is hard to fix and should be better awared by current LLM developers and researchers.
Limitations and Future Works. Currently, our benchmark is Chinese-only due to the language limitation of human annotators and moderators, which may cause an edge on performance for LLMs by Chinese companies such as Ernie-4.5 and Hunyuan Turbo-S (note the main message, significant performance decrease after modification, is not affected). Though our message to convey is already strong with the current results (and preliminary English translation tests in this paper suggest that LLMs will other struggle on the other languages), to expand such benchmark to multiple languages will be an important future work. A more important and fundamental avenue for future research is to find an effective way for LLMs to overcome the problem of recitation over reasoning without over-reliance on user’s clarifications or being too harsh on typos.
@@ Acknowledgements
The appendix has two parts. In Sec. A, we introduce the prompts used in our experiments; in Sec. B, we show more experiment results on our RoR-Bench.
Appendix A Prompts
A.1 Prompts and OpenAI-o1-1217 Response in Fig. 1
Due to space limit, we simplified the response of OpenAI-o1-1217 in Fig. 1. Here we show the full prompts and responses below:
[…]
See https://chatgpt.com/share/67edc40f-e658-800d-9bcd-7cb268a6f8c9 for another example of using the word “floating” and https://chatgpt.com/share/67ee1ba2-089c-800d-ad01-d7ee50dab39d for “drifting”. Interestingly, o1 can recognize the issue when we use the two words together (https://chatgpt.com/share/67ee1bda-5060-800d-ad44-b6c4e5741185), but not separately.
A.2 Prompts for the Judge
We use the same prompt for GPT-4o-1120 judges for all our experiments. We use Chinese prompts as the problems are Chinese; we provide English translations as well in the boxes below. Interestingly, we found the emphasis of “do not question the ground truth answer” very important; the recitation issue of LLMs is so severe that without such prompt, the LLM judge will question the problem instead of the student’s answer even with ground truth answer available.
# Chinese Prompt{CJK} UTF8gkai 你是一个专业的老师,现在有一道题目,你需要判断学生的回答是否和标准答案一致。题目和答案保证是绝对正确的,不会有错字,因此你要严格按照题目的字面意思评判。不要质疑标准答案有问题。 如果学生的回答和标准答案一致,请打1分;否则请打0分。标准答案经常只含有答案,而学生的回答一般会带有过程;你只需要根据学生的结论是否与标准答案相符做出评价。学生的结论和标准答案必须本质一样,但表述可以稍有不同;例如,学生的答案是50又1/2或50.5,标准答案是101/2,则也可以算作正确。 你的判断应遵循如下格式:你的输出在【评价】后开始。首先是对学生回答的分析(不超过300字),然后以“分数:[[0]]”或“分数:[[1]]”结束。 下面开始判断: 【题目】 (Problem description){CJK} UTF8gkai 【标准答案】 (Answer){CJK} UTF8gkai 【学生回答】 (LLM output)
@@ A.2 Prompts for the Judge
A.3 Prompts for Few-Shot In-Context Learning
@@
Please answer the following problems. The problems are guaranteed to be correct; please strictly follow the literal meaning of the problem. Here is an example: [Problem] Four people need to swim across a river. They have only one swimming ring, and at most two people can use it at the same time. Someone must carry the swim ring whenever it is used. The time it takes for each person to swim across the river individually is 1 minute, 2 minutes, 5 minutes, and 10 minutes respectively. If two people use the swim ring together to cross the river, the time it takes is equal to the time of the slower swimmer. The question is: how can all four people cross the river within 17 minutes? [Answer] Let the 1-minute and 2-minute people use the swim ring to cross the river first, which takes 2 minutes. Then the 1-minute person brings the swim ring back, taking 1 minute. Next, the 5-minute and 10-minute people cross the river together using the swim ring, which takes 10 minutes. After that, the 2-minute person brings the swim ring back, taking 2 minutes. Finally, the 1-minute and 2-minute people cross the river together again using the swim ring, taking 2 minutes. The total time spent is: 2 + 1 + 10 + 2 + 2 = 17 minutes. Now here is the problem you need to answer: [Problem] Four people need to swim across a river. They have only one swimming ring, and at most two people can use it at the same time. Anyone who cannot swim must use the swim ring, and it must be carried by someone while in use. The times it takes for each person to swim across the river individually are 1 minute, 2 minutes, 5 minutes, and 10 minutes respectively. Among them, the first three people can swim. If two people use the swim ring together to cross the river, the time required is equal to the time it takes for the slower person to cross the river alone. The question is: how can all four people cross the river within 17 minutes?
@@
Interestingly, when we test this English translation with OpenAI-o1-1217, we found o1, even with 1-shot, is again tricked into the classic paradigm that the swimming ring must be carried back. The ground truth answer of this target problem, however, is to directly let the third and fourth people use the swimming ring, and the first two people swim through the river, such that everything can be done within 10 minutes; no swimming ring needs to be taken back.
Appendix B More Experiment Results
B.1 Best-of-5 Results

### v3

@@ 4.2 Vision-based Problems
Evaluations. We evaluate $15$ cutting-edge VLMs, which are: GPT-4.5-Preview, OpenAI-o1-1217, GPT-4o-1120, Gemini-2.0 Pro-0205, GPT-4o-mini-0718, Gemini-2.0 Flash-0121, Qwen-2.5-VL-max, GLM-4v-Plus, Qwen-2.5-VL-72B, Claude 3.5 Sonnet, StepFun-1v-32k, Nova-Pro, Claude 3.7 Sonnet, SenseChat-Vision [60], and Qwen2.5-VL-7B. Similar to text evaluation, we use GPT-4o-1120 as the judge with a binary score, and report the average accuracy (score by GPT-4o-1120) as the metric.
@@
Table 2: Results on vision-based problems of RoR-Bench. All scores are binary and averaged over $5$ trials and $57$ problems, normalized to 01000-100 (higher is better). Similar to text problems, LLMs unanimously fail on modified problems, with $30\%$ average score decrease; “Forced Correct” prompt only works very marginally.
Results. Tab. 2 shows the result for all VLMs tested on RoR-Bench, which exhibits a $35\%$ perfomance decrease on average from original problems to the modified problems. Interestingly, we find GPT-4o-1120, GPT-4.5-Preview and OpenAI-o1-1217 to be significantly better on original problems, but much worse on modified problems; upon checking responses, we find that the OpenAI models listed above are much more likely to summarize the origin of the images, as we collect them usually from illustrations of famous visual effects (e.g. Mach bands and checker-shadow illusions). On the contrary, models like Claude 3.5 Sonnet and Claude 3.7 Sonnet usually do not explicitly summarize such visual effects. Such result indicates that 1) OpenAI models may be overfitting to usual test cases, and more importantly, 2) explicit summarization or knowledge retrieval, which already becomes a common practice for prompt-engineering [38, 86], is a double-edged sword; while they improve the performance on usual test cases, it may increase the risk of missing key details in the problem during summarization.
@@ 4.3 Is Few-Shot In-Context Learning the Cure?
Evaluations. We evaluate the same set of LLMs in Sec. 4.1  [^6] For better readability, we only show the most representative models in the main paper; see Appendix C.6 for details. For case 1 (adding original problems) mentioned above, we add a simple prompt mentioning the original problem and solution are an example (See Appendix B.2 for details). For case 2 (adding modified problems), we uniformly randomly select modified problems other than the current problem as shots; we test both 1-shot and 5-shot scenario.
Results. The results of the most represenatitive LLMs for few-shot In-Context Learning (ICL) is listed in Tab. 3; see Appendix C.6 for other LLMs. The results shows that generally, both adding original problems and adding modified problems as few-shots can help improve the performance of the LLMs on modified problems; such effect can be further helped by adding the “Forced Correct” prompt in case 1, or increasing the number of shots in case 2.
Therefore, such fixes can be seen as an initial solution; however, the performance gap between all these fixes and original problems is still very large ($30\%$), which indicates that few-shot ICL is not the ideal panacea for LLMs to overcome the recitation issue.
@@
Table 3: The results of adding original problems as 1-shot (case 1) or adding other modified problems as few-shot (case 2) sorted by average score on modified problems in our benchmark. Claude 3.7 Sonnet and Gemini-2.0 Flash-0121 are without long CoT (same for Tab. 4). Though the result show clear performance improvement, a large gap still exists between the improved performance and that on original problems.
4.4 Overconfidence in Solvability
As real-life problems can be ill-posed sometimes with no valid solution, a good LLM agent should possess the ability to discriminate such type of problems. However, as we examine the “no solution” problems in our benchmark (see Sec. 3.2 for details), we found that LLMs are particularly worse in correctly pointing out the problems with no solution, and often will make mistakes to make up a solution, as if injected by the mental seal that the problem is definitely solvable.
Evaluations. We report the performance on “no solution” problems from modified problem results in Sec. 4.1. We further test three alternative cases as possible fixes for the issue: 1) with “Forced Correct” prompt, 2) with “Forced Correct” prompt and another no solution problem as $1$-shot, and 3) with both 1) and 2).
@@ 4.4 Overconfidence in Solvability
Table 4: The scores for “no solution” problems and possible fixes, sorted by average score on such of problems. It is clearly shown that without any fixes, the average score for “no solution” problems is extremely low, showing the firm belief of LLMs that the given problem is solvable. While some LLMs, such as GPT-4.5-Preview, can be effectively corrected by adding “Forced Correct” (FC) prompts and other “no solution” problems as $1$-shot, other LLMs such as DeepSeek-R1 are still very stubborn.
Results. Tab. 4 shows the performance of the most representative LLMs on “no solution” problems as stated in Sec. 3.2 (see Appendix C.7 for other LLMs). Surprisingly, without any fixes, LLMs are unanimously stubborn on the belief that the given problem is solvable; not a single model achieves $15\%$ score on this type of problems. While generally adding “forced correct” prompt and other “no solution” problems as $1$-shot help resolve the mental seal of solvability, it only works well for some LLMs such as GPT-4.5-Preview, and is generally still far from satisfactory for most models.
Interestingly, DeepSeek-R1 struggles in recognizing unsolvable questions; we find that it has a firmer belief that the problem should be handled in the usual pattern even with directions that the problem should be taken literally, or several examples suggesting that it is more similar to a brain teaser; more often than other models, we found the form “the problem may suggest xxx, but maybe we should still consider xxx …” in its CoT. Overall, thinking models with long CoT are more likely to somewhat realize that the problem might contain traps, but assume the problem to be “normal” (i.e. closer to original), while normal LLMs often answer the problem unaware of the condition change. This suggests that long CoT might be a possible way to mitigate the issue, albeit not in the current status; more alignment is required to make its judgment on the problem more similar to humans.
4.5 Why Does Recitation Happen?
To address why recitation happens, we consider three possible reasons: 1) Dispersed attention, i.e., the model ignores the subtly changed condition due to insufficient attention weights; 2) Over-alignment and bad instruction following ability, i.e., the models stick to the “common” user intention and are reluctant to follow the problem in literal even with the “force correct prompt”, and 3) Solution paradigm overfitting, i.e., the model does not see the subtly changed problems in its training and thus performs poorly on out-of-distribution data.
To test the first reason, we add irrelevant text (The Thousand Character Classic/ 《千字文》, a Chinese poem) in front of each problem. We report the accuracy percentage change in Tab. 5:
[…]
DeepSeek-R1 -1.9 -2.9
Claude-3.5-Sonnet +1.0 +5.7
Gemini-2.0 Pro +1.0 +6.7
GPT-4o-mini-0718 -3.8 +1.9
GPT-4o-1120 -1.9 -1.9
Gemini-2.0 Flash +2.9 0
[…]
The result shows no or very slight performance change on modified problems on average; thus, attention dispersion is likely not the culprit.
For the second reason, we witness some cases (12 for o1 on text-based problems) where the model keeps adhering to the “usual” condition as in Appendix A.1 and Sec. 4.3. However, there are still many cases where o1 and DeepSeek-R1 fall into trap on our modified problems without noticing the condition change, with non-thinking models usually (if not always) ignoring them. Also, by comparing Tab. 1 and “case 1” in Tab. 3, the performance increase by adding “forced correct” prompt (instruction prompt) is roughly the same as adding the original problem (non-instruction prompt) as 1-shot (5.06% vs. 5.16% on average) which can stack (12.52% combined). Thus, overalignment/bad instruction following is partly the reason, but cannot account for most of the performance gap.
In conclusion, our hypothesis is that solution paradigm overfitting is the main culprit, while over-alignment / instruction following ability is also a factor. Long CoT gives the model more diverse reasoning paradigms and chances of self-reflection to mitigate the issue, but the model still needs to prevent overalignment to gain performance.
@@ 5 Discussion and Conclusion
In this work, we propose RoR-Bench, a multimodal Chinese benchmark which clearly reveals an alarming issue that current cutting-edge LLMs are unable to address even simple reasoning problems with conditions subtly shifted. Such phenomenon proved that LLMs are conducting recitation instead of reasoning when confronting seemingly classic problems. We found such issue can lead to dramatic performance loss ($50\%$) and is unable to be addressed by simple fixes such as adding instruction prompts or few-shots, indicating that such issue is hard to fix and should be better awared by current LLM developers and researchers.
Limitations and Future Works. Currently, our benchmark is Chinese-only due to the language limitation of human annotators and moderators, which may cause an edge on performance for LLMs by Chinese companies such as Ernie-4.5 and Hunyuan Turbo-S (note the main message, significant performance decrease after modification, is not affected). Though our message to convey is already strong with the current results (and preliminary English translation tests in this paper suggest that LLMs will other struggle on the other languages), to expand such benchmark to multiple languages will be an important but challenging future work (see Appendix A for detailed discussion). A more important and fundamental avenue for future research is to find an effective way for LLMs to overcome the problem of recitation over reasoning without over-reliance on user’s clarifications or being too harsh on typos.
@@ Acknowledgements
The appendix is organized as follows. In Sec. A, we discuss the possible challenges of expanding this benchmark to other languages. In Sec. B, we introduce the prompts used in our experiments, and in Sec. C, we show more experiment results on our RoR-Bench.
Appendix A Challenges of Expanding to Other Languages
One limitation of our benchmark is that it is Chinese-only in its current form. From our results, we tend to believe that the recitation over reasoning problem is language-agnostic. However, models encounter more severe phenomena with Chinese due to the potentially large meaning variance with minimal change (e.g. a single character); to extend to other languages such as English could be potentially very challenging. Here are some examples:
Example 1: In our benchmrk ,the original question is “…船从甲地开往乙地….”, and the modified question is “船从甲地漂往乙地”. Here, the word 开 strongly indicates that the engine is involved, while the word “漂” strongly indicates that the engine is not involved in Chinese. O1 translates “漂” as “floating” or “drifting”; however, we received feedback from English native speakers that such words are ambiguous in English with respect to the use of engine.
Example 2: the original question is “…相向而行…”, while the modified question is “相背而行”. Here, the former means “travel towards each other”, and the latter means “travel away from each other”. O1 gives a translation of “travel in the opposite directions”. However, “travel in the opposite directions” is ambiguous as traveling away from and traveling towards each other are both “traveling in the opposite directions”.
Thus, translating our benchmark into English is a non-trivial effort which requires extended efforts from native speakers; currently, we treat this as an important future work.
Appendix B Prompts
B.1 Prompts for the Judge
We use the same prompt for GPT-4o-1120 judges for all our experiments. We use Chinese prompts as the problems are Chinese; we provide English translations as well in the boxes below. Interestingly, we found the emphasis of “do not question the ground truth answer” very important; the recitation issue of LLMs is so severe that without such prompt, the LLM judge will question the problem instead of the student’s answer even with ground truth answer available. In our experiment, we find that the response of GPT-4o judge is usually in the following form: “the student’s answer is xxx, the ground truth is xxx. As they are different in xxx / essentially the same, the final score is 0 / 1.”
# Chinese Prompt 你是一个专业的老师,现在有一道题目,你需要判断学生的回答是否和标准答案一致。题目和答案保证是绝对正确的,不会有错字,因此你要严格按照题目的字面意思评判。不要质疑标准答案有问题。 如果学生的回答和标准答案一致,请打1分;否则请打0分。标准答案经常只含有答案,而学生的回答一般会带有过程;你只需要根据学生的结论是否与标准答案相符做出评价。学生的结论和标准答案必须本质一样,但表述可以稍有不同;例如,学生的答案是50又1/2或50.5,标准答案是101/2,则也可以算作正确。 你的判断应遵循如下格式:你的输出在【评价】后开始。首先是对学生回答的分析(不超过300字),然后以“分数:[[0]]”或“分数:[[1]]”结束。 下面开始判断: 【题目】(Problem description) 【标准答案】 (Answer) 【学生回答】 (LLM output)
@@ B.1 Prompts for the Judge
B.2 Prompts for Few-Shot In-Context Learning
@@
请回答下面的问题。题目保证没有错误,请严格按照题目字面所写的问题回答。以下是一个例子: 【问题】有四个人要在游过一条河,他们只有一个游泳圈,且每次最多只能两个人一起使用游泳圈游过河,使用游泳圈时必须有人携带 。四个人单独游过河的时间分别是 1 分钟、2 分钟、5 分钟、10 分钟。如果两人一起使用游泳圈游过河,所需要的时间就是游得慢的那个人单独游过河的时间。请问,他们如何在 17 分钟内全部游过河? 【答案】让 1 分钟和 2 分钟的人先一起使用游泳圈游过河,花费 2 分钟,然后 1 分钟的人带着游泳圈游回来,花费 1 分钟。 5 分钟和 10 分钟的人一起使用游泳圈游过河,花费 10 分钟,接着 2 分钟的人带着游泳圈游回来,花费 2 分钟。 1 分钟和 2 分钟的人再次一起使用游泳圈游过河,花费 2 分钟。 总共花费的时间为:2+1+10+2+2 = 17 分钟。 下面是你要回答的问题: 【问题】有四个人要在游过一条河,他们只有一个游泳圈,且每次最多只能两个人一起使用游泳圈游过河,不会游泳的人必须使用游泳圈,使用时必须有人携带 。四个人单独游过河的时间分别是 1 分钟、2 分钟、5 分钟、10 分钟。其中前三个人均会游泳。如果两人一起使用游泳圈游过河,所需要的时间就是游得慢的那个人单独游过河的时间。请问,他们如何在 17 分钟内全部游过河? Please answer the following problems. The problems are guaranteed to be correct; please strictly follow the literal meaning of the problem. Here is an example: [Problem] Four people need to swim across a river. They have only one swimming ring, and at most two people can use it at the same time. Someone must carry the swim ring whenever it is used. The time it takes for each person to swim across the river individually is 1 minute, 2 minutes, 5 minutes, and 10 minutes respectively. If two people use the swim ring together to cross the river, the time it takes is equal to the time of the slower swimmer. The question is: how can all four people cross the river within 17 minutes? [Answer] Let the 1-minute and 2-minute people use the swim ring to cross the river first, which takes 2 minutes. Then the 1-minute person brings the swim ring back, taking 1 minute. Next, the 5-minute and 10-minute people cross the river together using the swim ring, which takes 10 minutes. After that, the 2-minute person brings the swim ring back, taking 2 minutes. Finally, the 1-minute and 2-minute people cross the river together again using the swim ring, taking 2 minutes. The total time spent is: 2 + 1 + 10 + 2 + 2 = 17 minutes. Now here is the problem you need to answer:
[Problem] Four people need to swim across a river. They have only one swimming ring, and at most two people can use it at the same time. Anyone who cannot swim must use the swim ring, and it must be carried by someone while in use. The times it takes for each person to swim across the river individually are 1 minute, 2 minutes, 5 minutes, and 10 minutes respectively. Among them, the first three people can swim. If two people use the swim ring together to cross the river, the time required is equal to the time it takes for the slower person to cross the river alone. The question is: how can all four people cross the river within 17 minutes?
@@
Interestingly, when we test this English translation with OpenAI-o1-1217, we found o1, even with 1-shot, is again tricked into the classic paradigm that the swimming ring must be carried back. o3-mini (https://chatgpt.com/share/67f60f89-e8c8-800d-b7c0-c0fcffaaad18), o3-mini-high (https://chatgpt.com/share/67f60f60-f0b0-800d-93cd-2e770dc7cbb5) and Gemini-2.5 Pro
(https://g.co/gemini/share/3ebe9a57c6ff) all fell for the trap for 0-shot). The ground truth answer of this target problem, however, is to directly let the third and fourth people use the swimming ring, and the first two people swim through the river, such that everything can be done within 10 minutes; no swimming ring needs to be taken back.
Appendix C More Experiment Results
C.1 Results on Verifiable Subset of RoR-Bench
To further verify our findings without possible bias by LLM judge, we test several models on a manually picked verifiable subset of our benchmark. The result is listed in Tab. 6 and 7, which shows similar conclusions to our main paper.
Model Original Modified Modified+FC
DeepSeek-R1 92.2 37.4 34.8
DeepSeek-V3 80.4 23.9 24.8
Claude-3.5-Sonnet 79.6 34.8 37.8
Claude-3.7-Sonnet 78.7 39.6 41.3
Gemini-2.0 Flash 78.3 20.9 20.4
Mistral-Large-2 72.6 30.0 34.4
Table 6: Results on text-based verifiable subset of RoR-Bench (FC=“Forced Correct” prompt).
Model Original Modified
GPT-4o 94.4 20.4
Qwen-VL-max 80.0 42.2
Gemini-2.0 Flash 75.2 38.5
SenseChat-Vision 76.3 38.5
Qwen2.5-VL-72B-Instruct 64.1 43.7
GPT-4o-mini 61.1 49.3
StepFun-1v 54.4 52.2
Table 7: Results on vision-based verifiable subset of RoR-Bench (FC=“Forced Correct” prompt).
C.2 English Version of Fig. 1
The response for OpenAI-o3 of the problem in Fig. 1 can be seen at https://chatgpt.com/share/687d4e38-680c-800d-9578-442c6819a5d7 and https://chatgpt.com/share/687d4e0c-0444-800d-983c-c63067a67820. For Gemini-2.5 Pro, the response can be seen at https://g.co/gemini/share/01fcaeb71e18 and https://g.co/gemini/share/d8dceea41f17. While they sometimes realize the possible ambiguity in the problem (as shown in the second response for OpenAI-o3), they can still often solve the problem directly without noticing the subtle change.
C.3 English Version of Fig. 2
The response for OpenAI-o3 for our English translation of the modified text problems in Fig. 2 can be seen in https://chatgpt.com/share/687d3f20-76c8-800d-bf9d-8edd8fcc07e5, and Gemini-2.5 Pro in https://g.co/gemini/share/9ead78631a31. For the modified image problem, the response for OpenAI-o3 can be see in https://chatgpt.com/share/687d3fe2-dec4-800d-a748-e6ebdb087d48 (o3 answers no, but the reasoning is about neighbors, which is irrelevant; there is no illusion within each diamond), and Gemini-2.5 Pro in https://g.co/gemini/share/6bd5bff68c2d.
C.4 Best-of-5 Results